# What if Othello-Playing Language Models Could See?

### v1

@@ Abstract
Language models are often said to face a symbol grounding problem. While some argue that world understanding can emerge from text alone, others suggest grounded learning is more efficient. We explore this through Othello, where the board state defines a simplified, rule-based world. Building on prior work, we introduce VisOthello, a multi-modal model trained on move histories and board images. Using next-move prediction, we compare it to mono-modal baselines and test robustness to semantically irrelevant perturbations. We find that multi-modal training improves both performance and the robustness of internal representations. These results suggest that grounding language in visual input helps models infer structured world representations.
Machine Learning, ICML
@@ 1 Introduction
Does a language model truly understand what cat refers to? While no one fully grasps the essence of a cat, human language users know enough to use the word appropriately—recognizing cats in images, inferring them from descriptions, and using the term naturally in context. Whether mono-modal language models can achieve this level of grounding remains an open question.
Rather than debating whether symbol grounding is in principle impossible for mono-modal models (Mitchell & Krakauer, 2023; Mollo & Millière, 2023), this paper focuses on a more pragmatic hypothesis: that incorporating multiple modalities enables more efficient learning. This question is orthogonal to, but compatible with, the idea that mono-modal models can still induce a form of referential semantics (Søgaard, 2023; Huh et al., 2024).
To explore this hypothesis, we turn to the concept of world model—a system that learns to represent the latent state of an environment and predict its evolution over time. While many world model studies use high-dimensional or continuous environments (e.g., robotics or simulated physics), we use the game of Othello as a minimal, symbolic world with deterministic transitions and well-defined rules.
Prior work has used this setup to investigate emergent world representations by training and probing models—from small transformers (Li et al., 2023; Nanda et al., 2023) to large language models (LLMs) (Yuan & Søgaard, 2025)—on sequences of Othello moves. These studies provide evidence that language models can learn to track the board state, forming a rudimentary world model when trained on large-scale sequential data, though this learning remains confined to the text-only domain.
@@
We extend the Othello learning task to a multi-modal setting by introducing VisOthello, a multi-modal Othello model trained on sequences of move histories and their corresponding board images; see Figure 1. For each sequence of moves, we generate a corresponding sequence of board state images, with each image depicting the board at a specific time step. We then apply masking strategies to selected move tokens and train the model to predict the missing steps, using both the move history and associated visual context.
@@
To address Sub1, we compare VisOthello against several baselines on the task of next move prediction, where the model predicts the next token given a partial game sequence. We evaluate the performance across varying data scales to assess learning efficiency. For Sub2, we perform a semantically irrelevant perturbation analysis by rotating the board image during inference, assessing whether the models trained on original images remain robust and continue to predict legal moves accurately.
Our findings show that multi-modal training leads to better performance and faster learning than learning from text alone. In addition, multi-modal models exhibit greater robustness under semantically irrelevant perturbations, as demonstrated in our board rotation analysis. These results suggest that while language models can infer world state representations from text alone, grounding them with visual input enables more efficient and robust construction of such representations.
@@ 2.1 Training Paradigm
Different from prior works that train Othello models in an autoregressive manner by predicting moves step-by-step, we adopt a BERT-style masked language modeling approach for training VisOthello. This avoids the computational overhead and complexity of autoregressive generation, enabling efficient bidirectional reasoning over static visual-text inputs without framing the task as video modeling (for a detailed explanation, see Appendix E). Specifically, we train VisOthello based on VisualBERT (Li et al., 2019).
@@ 2.2 Input Representation
Textual input. Following prior works (Li et al., 2023; Karvonen et al., 2024), we represent each game as a sequence of moves, where each move at time step $t$ is treated as a token, denoted as $m_{t}$. Our vocabulary consists of 64 unique tokens, corresponding to the 64 tiles on the board; see Fig 1. Image input. In addition to the textual input, we provide the model with a sequence of corresponding board images. As demonstrated in Figure 1, each image $b_{t}$ represents the board state after moves $m_{1}$, $m_{2}$, …, $m_{t-1}$, and serves as visual context for predicting the next move $m_{t}$. To extract visual features, considering the differences between Othello board images and object images in ImageNet (Russakovsky et al., 2015), we pretrain an Othello-specific image encoder using a ResNet-18 backbone (He et al., 2016) to extract visual features:
@@
Train Size 0 1k 3k 5k 10k 20k
Othello-GPT 07.34 19.61 70.89 80.00 82.56 81.46
BERT-S 20.74 90.86 90.81 90.79 91.78 93.09
ResNET-18-S 06.71 60.78 71.27 75.58 81.74 88.90
VisOthello-S 23.64 90.82 93.87 93.25 94.23 93.57
BERT-P 20.74 90.35 91.33 91.91 92.14 92.65
ResNET-18-P 06.71 71.28 84.57 88.08 91.02 92.23
VisOthello-P 43.30 91.42 92.15 92.89 93.39 94.03
Table 1: Legal move accuracy (%) for next move prediction across different data sizes. We highlight the best performing model of each training set size in bold. $-P$ indicates the model is pretrained, while $-S$ indicates it is trained from scratch.
@@ 2.3 VisOthello Training
We train the VisOthello model using two types of masked language modeling (MLM) strategies to enhance its ability to learn meaningful representations of both textual and visual game sequences. MLM enables the model to develop a deeper understanding of game dynamics. The overall training objective $\mathcal{L}_{\text{total}}$ is defined as the sum of the masked modeling and the text-image prediction loss as follows.
@@
To align with the next-move prediction setup used in Othello-GPT (Li et al., 2023), we additionally apply future token masking to the game sequence with a 20% probability. Given a textual move sequence of $m_{1}$, $m_{2}$, …, $m_{s}$, we randomly select a step $t$ ($1\leq t\leq s$) as the prediction target, and then mask all future tokens from $m_{t}$, which are $m_{t}$, $m_{t+1}$, …, $m_{s}$. To prevent information leakage, we also mask all the image tokens that contains the future move information, which is $v_{t+1}$, $v_{t+2}$, …, $v_{s}$. This setup reduces dependence on bidirectional context and fits better with the next-move prediction setup in the Othello.
@@
We also adapt the sentence-image prediction task in original VisualBERT training for the Othello task. For a given sequence of image tokens, we replace the corresponding move sequence to a random sequence at a chance of 50%. The model is trained to distinguish whether the text and image sequences are from the same game via binary classification. This helps to train the model better learn implicit alignments between language and vision.
@@ 3 Experiments
(a) Legal move acc. with 1K samples
(b) Legal move acc. with 10K samples
(c) Legal move acc. with 20K samples
Figure 3: Comparison of models’ performance with and without board rotation across different training dataset sizes. The results demonstrate that multi-modal models maintain better performance under rotation compared to purely visual models. $-P$ indicates the model is pretrained, while $-S$ indicates it is trained from scratch.
@@ 3.1 Experimental Setups
To better assess the impact of multi-modal learning, we include text-only baselines such as Othello-GPT, BERT and vision-only model ResNet for direct comparison. We collect a total of 25,657 real game records from the eOthello website [^1] https://www.eothello.com/, which serve as the textual sequence inputs for our dataset. We generate the corresponding board images based on the dataset and split the dataset into training (80%), validation (5%), and test (15%) sets. When evaluating VisOthello, we use legal move accuracy. It measures whether the predicted move $m_{t}$, given the move history $m_{1}$, $m_{2}$, …, $m_{t-1}$, is valid under Othello’s rules. To assess the learning efficiency of mono-modal and multi-modal models, we train all models on the full dataset (20k samples) as well as on randomly sampled subsets of 1k, 3k, 5k, and 10k examples. For VisOthello, we use the best performing image encoder, ResNet-18 pretrained and fine-tuned on the full 20k dataset, for feature extraction. We also investigate the impact of pretraining by training each model either from scratch or from publicly available pretrained weights. Details see Appendix F and 2.
3.2 Experimental Results
Table 1 report the next legal move prediction accuracy of various models across different dataset sizes. Several key observations emerge from these results. Multi-modal learning is more sample efficient. VisOthello achieves high accuracy (over 91% in next legal move prediction) with as few as 1k training examples, while uni-modal models either require more data or fail to reach the same performance ceiling. This suggests that multi-modal learning is more sample-efficient—VisOthello shows stronger performance at smaller scales than uni-modal baselines. This observation aligns with previous work from Zhuang et al. (2024). Pretraining information is not consistently helpful. While pretraining improves performance in some cases, especially in low-data regimes, its effect is not consistent across modalities. For ResNet-18, pretraining provides a substantial boost at smaller dataset sizes, but the improvements diminish as more data becomes available. In contrast, for both BERT and VisOthello, pretraining does not consistently lead to significant gains across training sizes. This observation is consistent with prior findings by Yuan & Søgaard (2025), suggesting linguistic pretraining may offer limited benefit for rule-based environments such as Othello.
3.3 Probing Internal Representations
Results. Figure 2 shows F1 scores from linear probes trained to predict tile-level board states from selected layers of BERT and VisOthello. When models are randomly initialized (0 examples), VisOthello already encodes more board-relevant structure than BERT, achieving substantially higher probe performance in early layers. This may be attributed to the use of a ResNet encoder pretrained on Othello board images, which already encodes useful spatial structure. As training dataset increases, both models improve, but VisOthello consistently achieves higher scores—especially in deeper layers. After training on 20k examples, VisOthello reaches 77.55 F1 at Layer 18, compared to 62.28 for BERT. This suggests that VisOthello learns more accurate internal representations of the board state, benefiting from both multi-modal input and architectural modifications.
4 Semantically Irrelevant Perturbation
We focus on board rotation as a concrete instance, which each test board is rotated 180 degrees. As illustrated in Figure 4, this transformation corresponds to a spatial inversion: for image-based models, this involves rotating the game board in the input image; for language-based models, it requires flipping the row and column indices of the move representation (e.g., D3 becomes E6, resulting in a different move token ID). We apply the rotation only at test time, evaluating models that were trained on the original (unrotated) training data. All models are assessed on their ability to predict the next legal move, as described in Section 3. This allows us to examine whether models rely on absolute visual or positional cues, or whether they have learned more abstract, generalizable representations of the board state.
As shown in Figure 3, BERT remains robust under board rotation, maintaining 90–93% accuracy across settings. This is expected, as symbolic move sequences can be deterministically remapped (e.g., D3 to E6). In contrast, ResNet-18’s accuracy drops sharply to 28–35%, indicating a failure to learn rotation-invariant representations and a reliance on absolute spatial patterns. Without access to move history or turn information, it depends on ambiguous visual cues easily disrupted by rotation—a key limitation of purely visual models in tasks like Othello. VisOthello, combining ResNet’s visual features with BERT’s move encoding, sustains high accuracy (91–93%) even after rotation. Sequence information, including player turns and prior moves, helps disambiguate visual input, while language guidance stabilizes predictions under spatial transformations. This highlights the advantage of multi-modal grounding: by aligning perceptual input with symbolic context, VisOthello overcomes the spatial brittleness of vision-only models.
We then perform feature alignment in Appendix J.
5 Conclusion
We address the task of learning to play Othello and extend it to a multi-modal setting by introducing a novel model, VisOthello. Specifically, we examine whether access to visual state information improves sample efficiency and accelerates learning by comparing VisOthello to existing mono-modal baselines on next move prediction. To further assess the benefits of multi-modal grounding, we introduce a board rotation perturbation and conduct feature alignment analysis to evaluate whether the models learn more robust and aligned representations. Our findings suggest that grounding language models with visual input leads to more efficient and stable learning, validating the world model theory (detailed discussion see Appendix A and B).
Appendix A Potential Impacts
Appendix B Limitations
Moreover, we do not include comparisons with large-scale text-only language models, as these have been thoroughly investigated in prior work (Yuan & Søgaard, 2025). Given that pretraining on language alone does not necessarily enhance understanding of the structured reasoning inherent in Othello, scaling up to such models and benchmarking against them is not currently a priority. Instead, our use of lightweight language models offers a practical and efficient probe into how much language pretraining contributes to this domain.
Appendix C Ethics Statement
Appendix D Related Work
D.1 LLMs for Game Sequence Modeling
Using AI models to play games is not a new concept. Early models, such as AlphaGo, were designed to master gameplay by using predefined game rules and structured environments (Silver et al., 2016, 2017; Feng et al., 2023). Recently, modeling games with LLMs and examining their understanding of game dynamics has become a popular research direction in LLM cognitive probing. Li et al. (2023) train GPT-2 on synthetically generated Othello games, then use probing techniques to determine whether the model develops internal representations of the game state—effectively inferring a world model. Building on this work, Nanda et al. (2023) demonstrate that game-related knowledge is linearly encoded within the model. Following this line, research has expanded the scope of world knowledge acquisition in other scenarios with more advanced probing methods (Hao et al., 2023; Yun et al., 2023; Vafa et al., 2024). For instance, works train similar models with other game datasets, such as chess, maze and checkers, finding that the same encoding patterns hold in these more complex games (Karvonen, 2024; Spies et al., 2024; Joshi et al., 2024; Karvonen et al., 2024). More relevant to our work, Yuan & Søgaard (2025) extend the study beyond GPT-2, evaluating state-of-the-art LLMs (e.g., LLaMA-2 (Touvron et al., 2023), Qwen (Bai et al., 2023)) to assess their capacity for structured game knowledge representation. Hua et al. (2024) explore this phenomenon in multilingual settings, examining how language models encode and transfer game-related knowledge across different languages. Our work is the first to incorporate visual information in Othello game understanding, providing deeper insights into board state representations.
D.2 Multi-modal Alignment
A growing body of research explores cross-modal alignment as a lens to understand the extent to which language models can internalize and generalize knowledge from text-only inputs (Pereira et al., 2018; Caucheteux et al., 2022; Li et al., 2024a; Ngo & Kim, 2024). Notably, Merullo et al. (2023) demonstrate that visual representations can be effectively projected into the linguistic embedding space using simple linear transformations, revealing a surprising degree of structural compatibility between visual and textual modalities. Building on the theme, Li et al. (2024b) and Huh et al. (2024) argue that as model capacity increases, representations across modalities tend to converge toward a shared, modality-agnostic statistical structure of the world. Unlike prior work focused on aligning visual and linguistic representations of concrete objects, we extend this to abstract game mechanics, enabling deeper insight into how models understand structured environments from text alone.
Appendix E Model Design Motivation
We employ a masked language modeling (MLM) objective, following VisualBERT, rather than an autoregressive GPT-style objective, primarily to reduce computational cost and model complexity. Autoregressive models require sequential token generation and unidirectional attention, which becomes inefficient when integrating visual features: inserting images into the token stream introduces a pseudo-temporal dimension, effectively treating the input as a video sequence. This not only increases memory and compute requirements but also complicates the training dynamics without offering clear benefits for static visual-text reasoning tasks such as Othello. In contrast, MLM supports bidirectional attention and parallel training, allowing the model to leverage both visual and textual context efficiently. By avoiding unnecessary autoregressive structure, we preserve modeling capacity while significantly reducing resource demands.
Appendix F Baselines
@@
Appendix G Model Training Details
All models are trained for up to 1000 epochs, with validation performed every 10 epochs. We apply early stopping with a patience of 5 validation steps, and retain the checkpoint with the highest validation accuracy for final evaluation. Training is conducted on a single NVIDIA A100-40GB GPU. BERT and VisOthello are trained with a batch size of 128 and a learning rate of 1e-4, while ResNet is trained with a batch size of 512 using the same learning rate. Details of the dataset is shown in Table 2.
@@ Appendix G Model Training Details
Appendix H Rotation Demonstration
We demonstrate a board rotation example in Figure 4 .
Appendix I Ablation Study
We perform ablations to assess the impact of the modified components in VisOthello relative to the original VisualBERT.
We assess whether fine-tuning a ResNet model for image feature extraction is necessary. Specifically, we compare it with simpler alternatives that do not involve task-specific training. The first approach (Pooling) performs a downsampling operation on the raw image of size $600\times 600\times 3$ using a spatial pooling mechanism, reducing it to a 1200-dimensional embedding. The second approach (Area) employs pixel area relation interpolation, where the image is partitioned into fixed-size patches, pixel values within each region are averaged, and the resulting resized image is flattened into a 1200-dimensional vector. We also evaluate a ResNet-18 encoder without any fine-tuning on Othello images.To test the role of future token masking, we ablate this component by training VisOthello without it (denoted as W/O FTM) and compare performance against the full model.

### v2

@@ Abstract
Language models are often said to face a symbol grounding problem. While some have argued the problem can be solved without resort to other modalities, many have speculated that grounded learning is more efficient. We explore this question in Othello, a simplified, rule-based world that offers a controlled and interpretable testbed for studying world understanding. Building on prior work, we introduce VisOthello, a multi-modal model trained jointly on move sequences and board images. Using the Othello rule understanding task, we examine whether multi-modal learning provides advantages over text-only approaches. We further evaluate robustness under semantically irrelevant perturbations and analyze the consistency of cross-modal alignment. Our results suggest that multi-modal training not only improves performance and robustness but also promotes convergence toward shared internal representations across different model architectures.
What if Othello-Playing Language Models Could See?
[…]
@@ 1 Introduction
Does a language model truly understand what cat refers to? Of course, none of us fully know what a cat is in an absolute sense, but human language users know enough to use the word appropriately. We can identify cats in images, infer that the furry, mouse-loving pet someone just described is likely a cat, and use the term in context with ease. Whether mono-modal language models can achieve this level of grounding remains an open question.
This paper does not aim to engage with the discussion over whether symbol grounding is in principle impossible for mono-modal language models (Mitchell and Krakauer, 2023; Mollo and Millière, 2023). Instead, we focus on the hypothesis that the inclusion of multiple modalities can facilitate more efficient learning. The question is orthogonal, but entirely consistent with the idea that mono-modal language models can induce (a form of) referential semantics (Søgaard, 2023; Huh et al., 2024).
To test this hypothesis, i.e., to what extent multi-modal language models are more sample-efficient, we turn to the task of learning to play Othello with language models, a domain that offers a well-defined, symbolic environment with clear rules and a compact action space, making it an ideal testbed Li et al. (2023); Hua et al. (2024). Prior work has used this setup to investigate emergent world representations, training models ranging from small-scale language models (Li et al., 2023) to large language models (LLMs) Yuan and Søgaard (2025) to predict the next move from prior moves, with performance evaluated by next legal move accuracy to assess rule learning. A probing classifier is trained to investigate the representations learned for intermediate game states (e.g., my move vs. your move) Nanda et al. (2023). Evidence suggests that language models can learn to track the board state, which potentially forms a rudimentary world model, when trained on large amounts of sequential data.
@@
While Li et al. (2023) showed that text-only models can develop emergent world representations in Othello, our work extends this framework into the multi-modal regime, by introducing VisOthello, an Othello model trained on sequences of move histories and their corresponding board images (see Figure 1). For each sequence of moves, we generate a corresponding sequence of board state images, with each image depicting the board at a specific time step. We then apply masking strategies to selected move tokens and train the model to predict the missing steps, using both the move history and associated visual context.
@@
Sub3
Do multi- and mono-modal models learn
aligned representations?
To address Sub1, we compare VisOthello against several baselines on the task of next move prediction, where the model predicts the next token given a partial game sequence. We evaluate the performance across varying data scales to assess learning efficiency. For Sub2, we perform a semantically irrelevant perturbation analysis by rotating the board image during inference, assessing whether the models trained on original images remain robust and continue to predict legal moves accurately. Regarding Sub3, inspired by Lample et al. (2017); Li et al. (2024b), we apply two feature alignment techniques to project intermediate representations from different models into a shared vector space and evaluate their similarity.
We show that multi-modal training improves performance and sample efficiency over text-only training. We also observe that multi-modal models exhibit greater robustness to board rotations. Furthermore, through a feature alignment analysis, we find that representational similarity between models increases with more training data—suggesting that, despite differences in architecture and modality, models can converge on shared internal representations.
Contributions.
We are the first to compare the learning curves of mono-modal and multi-modal language models on the task of Othello move prediction and their internal representation learning. We evaluate model robustness by testing invariance to semantically irrelevant perturbations (i.e., board rotations). Additionally, we further study grounding by aligning the internal representations of different models and modalities using supervised and unsupervised methods. Our code is available at https://github.com/shin-ee-chen/multimodal-othello.
2 Related Work
2.1 LLMs for Game Sequence Modeling
Using AI models to play games is not a new concept. Early models, such as AlphaGo, were designed to master gameplay by using predefined game rules and structured environments (Silver et al., 2016, 2017; Feng et al., 2023). Recently, modeling games with LLMs and examining their understanding of game dynamics has become a popular research direction in LLM cognitive probing. Li et al. (2023) train GPT-2 on synthetically generated Othello games, then use probing techniques to determine whether the model develops internal representations of the game state—effectively inferring a world model. Building on this work, Nanda et al. (2023) demonstrate that game-related knowledge is linearly encoded within the model. Following this line, research has expanded the scope of world knowledge acquisition in other scenarios with more advanced probing methods (Hao et al., 2023; Yun et al., 2023; Vafa et al., 2024). For instance, works train similar models with other game datasets, such as chess, maze and checkers, finding that the same encoding patterns hold in these more complex games (Karvonen, 2024; Spies et al., 2024; Joshi et al., 2024; Karvonen et al., 2024). More relevant to our work, Yuan and Søgaard (2025) extend the study beyond GPT-2, evaluating state-of-the-art LLMs (e.g., LLaMA-2 Touvron et al. (2023), Qwen Bai et al. (2023)) to assess their capacity for structured game knowledge representation. Hua et al. (2024) explore this phenomenon in multilingual settings, examining how language models encode and transfer game-related knowledge across different languages. Our work is the first to incorporate visual information in Othello game understanding, providing deeper insights into board state representations.
2.2 Multi-modal Alignment
A growing body of research explores cross-modal alignment as a lens to understand the extent to which language models can internalize and generalize knowledge from text-only inputs Pereira et al. (2018); Caucheteux et al. (2022); Li et al. (2024a); Ngo and Kim (2024). Notably, Merullo et al. (2023) demonstrate that visual representations can be effectively projected into the linguistic embedding space using simple linear transformations, revealing a surprising degree of structural compatibility between visual and textual modalities. Building on the theme, Li et al. (2024b) and Huh et al. (2024) argue that as model capacity increases, representations across modalities tend to converge toward a shared, modality-agnostic statistical structure of the world. Unlike prior work focused on aligning visual and linguistic representations of concrete objects, we extend this to abstract game mechanics, enabling deeper insight into how models understand structured environments from text alone.
@@ 3.1 Training Paradigm
Different from prior works that train Othello models in an autoregressive manner by predicting moves step-by-step, we adopt a BERT-style masked language modeling approach for training VisOthello. This avoids the computational overhead and complexity of autoregressive generation, enabling efficient bidirectional reasoning over static visual-text inputs without framing the task as video modeling (for a detailed explanation, see Appendix B). Specifically, we train VisOthello based on VisualBERT (Li et al., 2019).
@@ 3.2 Input Representation
Textual input.
Following prior works Li et al. (2023); Karvonen et al. (2024), we represent each game as a sequence of moves, where each move at time step $t$ is treated as a token, denoted as $m_{t}$. Our vocabulary consists of 64 unique tokens, corresponding to the 64 tiles on the board. For example, C4 and E6 correspond to the 27th and 45th token in the vocabulary, respectively.
Image input.
In addition to the textual input, we provide the model with a sequence of corresponding board images. As demonstrated in Figure 1, each image $b_{t}$ represents the board state after moves $m_{1}$, $m_{2}$, …, $m_{t-1}$, and serves as visual context for predicting the next move $m_{t}$. To extract visual features, considering the differences between Othello board images and object images in ImageNet Russakovsky et al. (2015), we pretrain an Othello-specific image encoder using a ResNet-18 backbone He et al. (2016). We then extract visual features using this encoder, resulting in a visual embedding:
@@ 3.3 VisOthello Training
We train the VisOthello model using two types of masked language modeling (MLM) strategies to enhance its ability to learn meaningful representations of both textual and visual game sequences. MLM enables the model to develop a deeper understanding of game dynamics.
@@
To align with the next-move prediction setup used in Othello-GPT Li et al. (2023), we additionally apply future token masking to the game sequence with a 20% probability. Given a textual move sequence of $m_{1}$, $m_{2}$, …, $m_{s}$, we randomly select a step $t$ ($1\leq t\leq s$) as the prediction target, and then mask all future tokens from $m_{t}$, which are $m_{t}$, $m_{t+1}$, …, $m_{s}$. To prevent information leakage, we also mask all the image tokens that contains the future move information, which is $v_{t+1}$, $v_{t+2}$, …, $v_{s}$. This task encourages the model to rely on previous move sequence rather than future information when predicting the next move. By masking all future tokens and images beyond a randomly selected time step, the model is trained to make predictions in a uni-directional manner, similar to autoregressive models. This setup reduces dependence on bidirectional context and fits better with the next-move prediction setup in the Othello game.
@@
We also adapt the sentence-image prediction task in original VisualBERT training for the Othello task. For a given sequence of image tokens, we replace the corresponding move sequence to a random sequence at a chance of 50%. The model is trained to distinguish whether the text and image sequences are from the same game via binary classification. This task helps to train the model better learn implicit alignments between language and vision.
Training objective.
The overall training objective $\mathcal{L}_{\text{total}}$ is defined as the sum of the masked modeling loss and the text-image prediction loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{mask}}+\mathcal{L}_{\text{ti}},$
where the masked modeling loss $\mathcal{L}_{\text{mask}}$ combines two components:
$\mathcal{L}_{\text{mask}}=\alpha\cdot\mathcal{L}_{\text{random}}+(1-\alpha)\cdot\mathcal{L}_{\text{future}}.$
Here, $\mathcal{L}_{\text{random}}$ is the random token masking loss, $\mathcal{L}_{\text{future}}$ is the future token masking loss, $\mathcal{L}_{\text{ti}}$ is the text-image prediction loss, and $\alpha$ is set to 0.8.
@@ 4 Experiments
In this section, we evaluate whether incorporating visual information improves learning efficiency and grounding in the Othello game setting.
@@ 4.1 Experimental Setups
Compared models.
To better assess the impact of multi-modal learning, we include text-only and vision-only baselines for direct comparison.
@@
Datasets.
We collect a total of 25,657 real game records from the eOthello website, [^1] https://www.eothello.com/ which serve as the textual sequence inputs for our dataset. For each recorded step in a game, we generate a corresponding image to capture its visual state. As summarized in Table 1, this process yields approximately 1.56 million images, averaging around 60.8 images per game across all splits. We split the dataset into training (80%), validation (5%), and test (15%) sets, while maintaining a consistent number of images per game across all splits.
@@
Train Size 0 1k 3k 5k 10k 20k
[…]
Table 2: Legal move accuracy (%) for next move prediction in different models across different data sizes. We report mean $\pm$ std over 3 runs, and highlight the best performing model of each training set size in bold. P-P indicates the model is pretrained, while S-S indicates it is trained from scratch.
[…]
Table 3: Exact match accuracy (%) for next move prediction in different models across different data sizes. For Othello-GPT, BERT, ResNet, and VisOthello we report mean $\pm$ std over 3 runs.
Evaluation metrics.
When evaluating VisOthello, we adopt the same setup as the future token masking objective following Li et al. (2023): to predict the legal move at step $m_{t}$, we mask all future tokens starting from $m_{t}$, and all image embeddings from $v_{t+1}$ onward, to prevent information leakage from future states. This setup assesses whether a model can learn the underlying rules of Othello game from sequential move data, in contrast to AlphaZero (Silver et al., 2017), which focus on the winning strategy. Accordingly, we use legal move accuracy as our evaluation metrics. Specifically, we evaluate whether the predicted move $m_{t}$, given the move history $m_{1}$, $m_{2}$, …, $m_{t-1}$, is valid under Othello’s rules, respectively. We also report exact match accuracy, which measures whether the predicted move $m_{t}$ exactly matches the ground-truth move, reflecting the model’s ability to replicate expert gameplay.
Training details.
To assess the learning efficiency of mono-modal and multi-modal models, we train all models on the full dataset (20k samples) as well as on randomly sampled subsets of 1k, 3k, 5k, and 10k examples. Training and evaluation are conducted with three random seeds (5, 12, and 42). For VisOthello, we use the best performing image encoder, ResNet-18 pretrained and fine-tuned on the full 20k dataset, for feature extraction. We also investigate the impact of pretraining by training each model either from scratch or from publicly available pretrained weights. [^2] For BERT, we use google-bert/bert-large-uncased; the pretrained ResNet-18 is from the HuggingFace transformers library; the pretrained VisualBERT weights are from the Volta framework. For training details, see Appendix C.
4.2 Experimental Results
Table 2 and 3 report the exact match and next legal move prediction accuracy of various models across different dataset sizes. Several key observations emerge from these results.
Multi-modal learning is more sample efficient.
VisOthello achieves high accuracy (over 91% in next legal move prediction) with as few as 1k training examples, while uni-modal models either require more data or fail to reach the same performance ceiling. This suggests that multi-modal learning is more sample-efficient—VisOthello shows stronger performance at smaller scales than uni-modal baselines. This observation aligns with previous work from Zhuang et al. (2024).
Pretraining information is not consistently helpful.
While pretraining improves performance in some cases, especially in low-data regimes, its effect is not consistent across modalities. With ResNet-18, pretraining is very helpful at small sizes, but its effect decreases as the dataset gets larger. In contrast, for both BERT and VisOthello, pretraining does not consistently lead to significant gains across training sizes. This observation is consistent with prior findings by Yuan and Søgaard (2025), which suggest that linguistic pretraining may offer limited benefit for structured, rule-based environments such as Othello.
4.3 Ablation Study
To evaluate the contribution of the modified components in VisOthello relative to the original VisualBERT, we conduct ablation studies focusing on the image encoder and the future token masking strategy.
Image encoder.
We test whether fine-tuning a ResNet model is necessary for extracting image features. For comparison, we consider three alternatives that do not involve task-specific adaptation: (a) a simple pooling projection that downsamples the raw $600\times 600\times 3$ image into a 1200-dimensional embedding, (b) an Area projection that partitions the image into patches, averages pixel values within each patch, and flattens the resized image into a 1200-dimensional vector, and (c) a ResNet-18 encoder without fine-tuning on Othello images.
Future token masking.
To test the necessity of future token masking (FTM), we train VisOthello without this component (denoted as W/O FTM) and compare performance with the full model.
Results.